# Real-Time Flying Object Detection with YOLOv8

### v1

@@ Abstract
This paper presents a generalized model for real-time detection of flying objects that can be used for transfer learning and further research, as well as a refined model that is ready for implementation. We achieve this by training our first (generalized) model on a data set containing 40 different classes of flying objects, forcing the model to extract abstract feature representations. We then perform transfer learning with these learned parameters on a data set more representative of “real world” environments (i.e. higher frequency of occlusion, small spatial sizes, rotations, etc.) to generate our refined model. Object detection of flying objects remains challenging due to large variance object spatial sizes/aspect ratios, rate of speed, occlusion, and clustered backgrounds. To address some of the presented challenges while simultaneously maximizing performance, we utilize the current state of the art single-shot detector, YOLOv8, in an attempt to find the best trade-off between inference speed and mAP. While YOLOv8 is being regarded as the new state-of-the-art [16], an official paper has not been provided. Thus, we provide an in depth explanation of the new architecture and functionality that YOLOv8 has adapted. Our final generalized model achieves an mAP50-95 of 0.685 and average inference speed on 1080p videos of 50 fps. Our final refined model maintains this inference speed and achieves an improved mAP50-95 of 0.835.
1 Introduction/Background/Motivation
@@
Numerous recent events have demonstrated the malicious use of drones. Over the past few months, there have been reports of assassination attempts via drones with small explosive payloads [17], drug deliveries to state prisons [19], and surveillance of U.S. Border Patrol by smugglers [5] to exploit weaknesses. While research indicates that drone usage is expected to increase exponentially [15], detection technology has yet to provide reliable and accurate results. Drones and mini UAVs present a stealth capability and can avoid detection by most modern radar systems due to their small electromagnetic signature. They are also small, highly maneuverable, and omit low levels of noise. This, along with the ease of access provides a natural incentive for drones to remain an integral part of modern warfare and illegal activities. While methods such as radio and acoustic detection have been proposed as solutions, they are currently known to be inaccurate [4]. This motivates the integration of a visual detector in any such detection system. The U.S. Border Patrol implements real-time object detection from digital towers to monitor people and motor vehicles [2], but are not currently known to implement drone detection. Drone detection in this environment is challenging due to the cluttered desert background and distance that drones survey from. [6]. The farther the drone is from cameras, the more difficult it will be to detect and classify it, as the object will convey less signal in the input space to the model.
@@
Our primary objective is to provide a generalized real-time flying object detection model that can be used by others for transfer learning or further research, as well as a refined model that is ready to use “out of the box” for implementation  [14]. We define a generalized model as one that has good detection and classification performance on a large number of classes at higher resolutions while maintaining a reasonable frame rate (1080p : 30-60 frames per second). Instead of just training our model on drones, we train on a data set containing 40 flying object categories to force the model to learn more abstract feature representations of flying objects. Then, we transfer learn these weights on a final data set containing more instances of “real world“ environments (i.e. higher frequency of occlusion, small spatial sizes, rotations, etc.). This in turn will lead to a more refined, ready to implement real-time flying object detection model. To maximize our model’s performance, we use the latest state-of-the-art single-shot detector, YOLOv8. Currently, single-stage detectors are the de-facto architecture choice for fast inference speeds. This choice comes at the expense of exchanging the higher accuracy you would typically expect from a two-state detector. While YOLOv8 is being regarded as the new state-of-the-art [16], an official paper has yet to be released. This motivates our secondary objective, which is to explain the new architecture and functionality that YOLOv8 has adapted.
@@ 2 Approach
We chose the YOLOv8 architecture under the assumption that it would provide us the highest probability of success given the task. YOLOv8 is assumed to be the new state-of-the-art due to its higher mAPs and lower inference speed on the COCO dataset. However, an official paper has yet to be released. It also specifically performs better at detecting aerial objects LABEL:fig:YOLOv8_average_mAP_against_cats. We implement the code from the Ultralytics repository. We decide to implement transfer learning and initialize our models with pre-trained weights to then begin training on the custom data set. These weights are from a model trained on the COCO dataset. Due to only having access to a single NVIDIA RTX 3080 and 3070, a greedy model selection/hyper-parameter tuning approach was chosen. We first train a version of the small, medium, and large versions of the model with default hyper-parameters for 100 epochs. Then, we decide which model is optimal for our use case given the trade off between inference speed and mAP-50-95 on the validation set. After the model size is selected, a greedy hyper-parameter search is conducted with 10 epochs per each set of hyper-parameters. The model with the optimal hyper-parameters trains for 163 epochs to generate the generalized model. After this model learns abstract feature representations for a wide-array of flying objects, we then transfer learn these weights to a data set that is more representative of the real world [1] to generate the refined model. This data set contains 3 classes: helicopter, plane, and drone, with very high variance in object spatial sizes. For evaluation, we are particularly interested in evaluating mAP50-95 and inference speed, as these are the most common measures of success across most object detection algorithms. Due to the large class imbalance, poor performance on the validation set was anticipated in the minority classes. However, this was not observed LABEL:fig:confusion_matrix.
@@
Mean average precision (mAP) is one of the most used evaluation metrics for object detection. mAP takes the average precision (AP) over all classes and computes them at a pre-specified IoU threshold. To define precision we need to define true positives and false positives for object detection. A true positive will be determined when the IoU between the predicted box and ground truth is greater than the set IoU threshold, while a false positive will have the IoU below that threshold. Then, precision can be defined as $\dfrac{tp}{tp+fp}$. We take the mean over a class by iterating over a set of thresholds and averaging them. For mAP50-95, we take steps of 0.05 starting from an IoU threshold of 0.5 and stopping at 0.95. The average precision over this interval is the class AP. Do this for all classes and take the average over them and we generate the mAP50-95.
2.1 Model Choice and Evaluation
@@
We evaluate small, medium, and large versions of the models to determine an optimal trade off between inference speed and mAP50-95 to then optimize the hyper-parameters. The small, medium, and large models have (11151080, 25879480, & 43660680) parameters and (225,295, & 365) layers respectively. After training the models we see there is a noticeable increase in mAP50-95 between small and medium models (0.05), but not much delta between medium and large (0.002). We also see that small, medium, and large infer at 4.1, 5.7, and 9.3 milliseconds respectively on the validation set. However, our original goal is to reach an average inference speed between 30 to 60 frames for 1080p. When testing the medium size model on multiple 1080p HD videos, we observe an average total speed (pre-proccess speed(0.5ms) + inference speed(17.25ms) + post-process speed(2ms)) of 19.75 ms (50 frames per second), which aligns with our primary objective. This leads to our selection of the medium size model to begin tuning hyper-parameters.
@@
Due to a lack of computational resources, we evaluate 10 epochs for each set of hyper-parameters as an indicator for the potential performance of additional epochs. We observe that this assumption is correct, as training with the optimal set of hyper-parameters achieves better performance at epoch 100 compared to default hyper-parameters (0.027)LABEL:fig:YOLOv8_mAP50-95_val We choose the best hyper-parameters based on validation mAP50-95 as batch size of 16, stochastic gradient descent (SGD) as the optimizer, momentum of 0.937, weight decay of 0.01, classification loss weight $\lambda_{cls}$ = 1, box loss weight $\lambda_{box}$ = 5.5, and distribution focal loss weight $\lambda_{dfl}$ = 2.5. After training for 163 epochs we achieve an mAP50-95 of 0.685 and an average inference speed on 1080p videos of 50 fps.
@@ 2.2 Loss Function and Update Rule
Where 1 is the generalized loss function incorporating the individual loss weights and a regularization term with weight decay $\phi$, 2 is the velocity term with momentum $\beta$, and 3 which is the weight update rule and $\eta$ is the learning rate. The specific YOLOv8 loss function can be defined as:
@@
  • $q_{(x,y)+/-1}$ are the nearest predicted boxes IoUs (left and right) $\in c^{*}_{x,y}$
@@
Each cell then determines its best candidate for predicting the bounding box of the object. This loss function includes the CIoU (complete IoU) loss proposed by Zheng et al.[22] as the box loss, the standard binary cross entropy for multi-label classification as the classification loss (allowing each cell to predict more than 1 class), and the distribution focal loss proposed by Li et al.[10] as the 3rd term.
Figure 1: YOLOv8 Validation
3 Experiments and Results
@@ 3.1 Model Confusion and Diagnosis
Figure 2: Feature Activation maps for the F-14 and F-18 fighter jets. From left to right we have the four stages of the model’s CSPDarkNet53 backbone
@@
Figure 3: (a) picks up the drone, (b) picks up tree top granularity - tree tops are more granular than stumps, (c) granular version of layer (b), (d) an outlier, texturized analysis of what the object is.
@@
One of the primary challenges in object detection is dealing with data sets with low inter-class variance, i.e., multiple classes that look similar to each other compared to the rest of the labels. Take, for example, the F-14 and F-18. Both have similar-looking wing shapes, two rudders, an engine, a cockpit, and a respective payload. In this confusion matrix 1, the model is most likely to mis-classify an F-14 as an F-18. This type of mis-classification typically affects classes in categories with low inter-class variance amongst themselves. Visualizing activation maps [21] is a technique that helps us understand what pixels in the input image are important for determining its class.
@@
Generally, deeper layers in CNNs extract more granular/complex/low level feature representations. YOLOv8 incorporates this idea into its architecture by having repeating modules and multiple detection heads when making its prediction. For our experimentation, we use MMYolo [21] to create activation maps at different stages of our backbone. We expect to some sense of differentiation in the different feature maps. If our model shows similar feature activations for F-14s and F-18s, we can say that may be the reason for class confusion.
@@
MMYolo [21] by Yamaguchi et al. is an open-source toolbox for YOLO series algorithms based on PYTorch. MMYolo can decompose the most popular YOLO algorithms, making them easily customizable and ready for analysis. For our analysis, we employed MMYolo to first convert the weights from .pt (Pytorch model) to .pth (State dictionary file, i.e., weights, bias, etc.) and second visualize the different activation maps of YOLOv8 during inference. MMYolo allows you to specify model type, weight file, target layer, and channel reduction.
@@
YOLOv8 6 uses CSPDarknet53 [13] as its backbone, a deep neural network that extracts features at multiple resolutions (scales) by progressively down-sampling the input image. The feature maps produced at different resolutions contain information about objects at different scales in the image and different levels of detail and abstraction. YOLOv8 can incorporate different feature maps at different scales to learn about object shapes and textures, which helps it achieve high accuracy in most object detection tasks. YOLOv8 backbone consists of four sections, each with a single convolution followed by a c2f module [16]. The c2f module is a new introduction to CSPDarknet53. The module comprises splits where one end goes through a bottleneck module(Two 3x3 convolutions with residual connections. The bottleneck module output is further split N times where N corresponds to the YOLOv8 model size. These splits are all finally concatenated and passed through one final convolution layer. This final layer is the layer we will get the activations.
@@
This figure 2 shows the original F-14 and F-18 images and the activations of the four c2f stages in the network, with each stage being more profound in the network from the second image right. The Activation Map corresponding to the shallowest c2f module shows the broadest activation. This module detects the two wings of the aircraft and determines that this object is a plane. The second activation map corresponds to the second c2f module in our backbone. It shows strong activations at different components of the aircraft, such as locating the wings, body, cockpit, and payload. It appears that this layer is attempting to infer what kind of aircraft is being presented in the image by highlighting these features. The third activation map is starting to dive into the individual textures of the components of the aircraft, presumably checking for minute differences in the jets structure. Finally, the model’s final c2f module activates extremely fine-grained details and outlines in the respective images. These similar feature activation maps could be the reason that the model confuses the two.
3.2 Model Examples and Results
To highlight our results, we address three challenging conditions: (1) detecting and classifying extremely small objects, (2) identifying flying objects that blend into their background, and (3) classifying different types of flying objects. We examined the performance of our generalized model,[11], against these challenges. This is demonstrated in Figure LABEL:fig:GeneralizedModel, which features four images that represent the bird, drone, passenger airplane, and V22 classes.
The first of the four images showcases the model’s ability to identify distant birds. In the second image, the model was put to the test against a very small drone that occupied only .026% of the image size while also blending in with its background. The model still resulted in the correct detection and classification of the drone. The third image shows the model’s ability to identify a minute passenger airplane of size 0.063% of the image, which is also blended into its surroundings. Finally, the fourth image features a V22 aircraft, which is an underrepresented class and accounts for only 3.57% of the entire dataset. A V22 can easily be mistaken as a drone due to its vertical propeller positioning. Despite these two characteristics and only taking up 0.14% of the entire image, the image exhibits the model’s ability to still identify the V22 with impressive accuracy, achieving a confidence score of 0.83.

### v2

@@ Abstract
This paper presents a generalized model for real-time detection of flying objects that can be used for transfer learning and further research, as well as a refined model that achieves state-of-the-art results for flying object detection. We achieve this by training our first (generalized) model on a data set containing 40 different classes of flying objects, forcing the model to extract abstract feature representations. We then perform transfer learning with these learned parameters on a data set more representative of “real world” environments (i.e. higher frequency of occlusion, very small spatial sizes, rotations, etc.) to generate our refined model. Object detection of flying objects remains challenging due to large variances of object spatial sizes/aspect ratios, rate of speed, occlusion, and clustered backgrounds. To address some of the presented challenges while simultaneously maximizing performance, we utilize the current state-of-the-art single-shot detector, YOLOv8, in an attempt to find the best trade-off between inference speed and mean average precision (mAP). While YOLOv8 is being regarded as the new state-of-the-art [19], an official paper has not been released as of yet. Thus, we provide an in-depth explanation of the new architecture and functionality that YOLOv8 has adapted. Our final generalized model achieves a mAP50 of 79.2%, mAP50-95 of 68.5%, and an average inference speed of 50 frames per second (fps) on 1080p videos. Our final refined model maintains this inference speed and achieves an improved mAP50 of 99.1% and mAP50-95 of 83.5%.
1 Introduction
@@
Numerous recent events have demonstrated the malicious use of drones. Over the past few months, there have been reports of assassination attempts via drones with small explosive payloads [20], drug deliveries to state prisons [22], and surveillance of the United States (U.S.) Border Patrol by smugglers [7] to exploit weaknesses. While research indicates that drone usage is expected to increase exponentially [17], detection technology has yet to provide reliable and accurate results. Drones and mini unmanned aerial vehicles (UAVs) present a stealth capability and can avoid detection by most modern radar systems due to their small electromagnetic signature. They are also small, highly maneuverable, and omit low levels of noise. This, along with the ease of access, provides a natural incentive for drones to remain an integral part of modern warfare and illegal activities. While methods such as radio and acoustic detection have been proposed as solutions, they are currently known to be inaccurate [6]. This motivates the integration of a visual detector in any such detection system. The U.S. Border Patrol implements real-time object detection from digital towers to monitor people and motor vehicles [4], but is not currently known to implement drone detection, which may explain the recent undetected illegal patrolling. Drone detection in this environment is challenging due to the cluttered desert background and the distance that drones survey from [8]. The farther the drone is from cameras, the more difficult it will be to detect and classify it, as the object will convey less signal in the input space to the model.
@@
Our primary objective is to provide a generalized real-time flying object detection model that can be used by others for transfer learning or further research, as well as a refined model that is ready to use “out of the box” for implementation  [16]. We define a generalized model as one that has good detection and classification performance on a large number of classes at higher resolutions while maintaining a reasonable frame rate (1080p : 30-60 frames per second). Instead of just training our model on drones, we train on a data set containing 40 different flying object categories to force the model to learn more abstract feature representations of flying objects. Then, we transfer learn these weights on a final data set containing more instances of “real world“ environments (i.e. higher frequency of occlusion, small spatial sizes, rotations, etc.). This in turn will lead to a more refined, ready-to-implement real-time flying object detection model. To maximize our model’s performance, we use the latest state-of-the-art single-shot detector, YOLOv8. Currently, single-stage detectors are the de-facto architecture choice for fast inference speeds. This choice comes at the expense of exchanging the higher accuracy you would typically expect from a two-state detector. While YOLOv8 is being regarded as the new state-of-the-art [19], an official paper has yet to be released. This motivates our secondary objective, which is to explain the new architecture and functionality that YOLOv8 has adapted.
2 Materials and Methods
@@
We chose the YOLOv8 architecture under the assumption that it would provide us with the highest probability of success given the task. YOLOv8 is assumed to be the new state-of-the-art due to its higher mean average precisions (mAPs) and lower inference speed on the COCO dataset. However, an official paper has yet to be released. It also specifically performs better at detecting aerial objects [Figure 9]. We implement the code for YOLOv8 from the Ultralytics repository. We decide to implement transfer learning and initialize our models with pre-trained weights to then begin training on the custom data set. These weights are from a model trained on the COCO dataset. Due to only having access to a single NVIDIA RTX 3080 and 3070, a greedy model selection/hyper-parameter tuning approach was chosen. We first train a version of the small, medium, and large versions of the model with default hyper-parameters for 100 epochs. Then, we decide which model is optimal for our use case given the trade-off between inference speed and mAP-50-95 on the validation set. After the model size is selected, a greedy hyper-parameter search is conducted with 10 epochs per each set of hyper-parameters. The model with the optimal hyper-parameters trains for 163 epochs to generate the generalized model. After this model learns abstract feature representations for a wide array of flying objects, we then transfer learn these weights to a data set that is more representative of the real world [1] to generate the refined model. This data set contains 3 classes: helicopter, plane, and drone, with very high variance in object spatial sizes. For evaluation, we are particularly interested in evaluating mAP50-95 and inference speed, as these are the most common measures of success across most object detection algorithms. Due to the large class imbalance, poor performance on the validation set was anticipated in the minority classes. However, this was not observed [Figure 2].
@@
Mean average precision (mAP) is one of the most used evaluation metrics for object detection. mAP takes the average precision (AP) over all classes and computes them at a pre-specified Intersection over Union (IoU) threshold. To define precision, we need to define true positives and false positives for object detection. A true positive will be determined when the IoU between the predicted box and ground truth is greater than the set IoU threshold, while a false positive will have the IoU below that threshold. Then, precision can be defined as $\dfrac{tp}{tp+fp}$. We take the mean over a class by iterating over a set of thresholds and averaging them. For mAP50-95, we take steps of 0.05 starting from an IoU threshold of 0.5 and stopping at 0.95. The average precision over this interval is the class AP. Do this for all classes and take the average over them and we generate the mAP50-95.
2.1 Generalized Model Choice and Performance
@@
We evaluate small, medium, and large versions of the models to determine an optimal trade-off between inference speed and mAP50-95 to then optimize the hyper-parameters. The small, medium, and large models have (11151080, 25879480, & 43660680) parameters and (225, 295, & 365) layers respectively. After training the models, we see there is a noticeable increase in mAP50-95 between small and medium models (0.05), but not much delta between medium and large (0.002). We also see that small, medium, and large infer at 4.1, 5.7, and 9.3 milliseconds respectively on the validation set. However, our original goal is to reach an average inference speed between 30 to 60 frames for 1080p. When testing the medium-size model on multiple 1080p HD videos, we observe an average total speed (pre-process speed (0.5ms) + inference speed (17.25ms) + post-process speed (2ms)) of 19.75 ms (50 frames per second), which aligns with our primary objective. This leads to our selection of the medium-size model to begin tuning hyper-parameters.
@@
Due to a lack of computational resources, we evaluate 10 epochs for each set of hyper-parameters as an indicator for the potential performance of additional epochs. We observe that this assumption is correct, as training with the optimal set of hyper-parameters achieves better performance at epoch 100 compared to default hyper-parameters (0.027) [Figure 2]. We choose the best hyper-parameters based on validation mAP50-95 as batch size of 16, stochastic gradient descent (SGD) as the optimizer, momentum of 0.937, weight decay of 0.01, classification loss weight $\lambda_{cls}$ = 1, box loss weight $\lambda_{box}$ = 5.5, and distribution focal loss weight $\lambda_{dfl}$ = 2.5. After training for 163 epochs, we achieve a mAP50-95 of 0.685 and an average inference speed on 1080p videos of 50 fps.
@@ 2.2 Loss Function and Update Rule
(1) is the generalized loss function incorporating the individual loss weights and a regularization term with weight decay $\phi$, (2) is the velocity term with momentum $\beta$, and (3) is the weight update rule with $\eta$ as the learning rate. The specific YOLOv8 loss function can be defined as:
@@
  • $q_{(x,y)+/-1}$ are the nearest predicted boxes IoUs (left and right) $\in c^{*}_{x,y}$.
@@
Each cell then determines its best candidate for predicting the bounding box of the object. This loss function includes the complete IoU (CIoU) loss proposed by Zheng et al. [25] as the box loss, the standard binary cross entropy for multi-label classification as the classification loss (allowing each cell to predict more than 1 class), and the distribution focal loss proposed by Li et al. [12] as the 3rd term.
Figure 1: Confusion matrix for all classes.
Figure 2: YOLOv8 validation mAP50-95.
@@ 2.3 Model Confusion and Diagnosis
Figure 3: Feature activation maps for the F-14 and F-18 fighter jets. From left to right, we have the four stages of the model’s CSPDarkNet53 backbone.
@@
Figure 4: From left to right, (1) picks up the drone, (2) picks up tree top granularity - tree tops are more granular than stumps, (3) granular version of layer (2), (4) an outlier, texturized analysis of what the object is.
@@
One of the primary challenges in object detection is dealing with data sets with low inter-class variance, i.e. multiple classes that look similar to each other compared to the rest of the labels. Take, for example, the F-14 and F-18, which are displayed in Figure 3. Both have similar-looking wing shapes, two rudders, an engine, a cockpit, and a respective payload. In this confusion matrix [Figure 2], the model is most likely to misclassify an F-14 as an F-18. This type of misclassification typically affects classes in categories with low inter-class variance amongst themselves. Visualizing activation maps [24] is a technique that helps us understand what pixels in the input image are important for determining its class.
@@
Generally, deeper layers in CNNs extract more granular/complex/low-level feature representations. YOLOv8 incorporates this idea into its architecture by having repeating modules and multiple detection heads when making its prediction. For our experimentation, we use MMYolo [24] to create activation maps at different stages of our backbone. We expect some sense of differentiation in the different feature maps. If our model shows similar feature activations for F-14s and F-18s, we can say that may be the reason for class confusion.
@@
MMYolo [24] by Yamaguchi et al. is an open-source toolbox for YOLO series algorithms based on PYTorch. MMYolo can decompose the most popular YOLO algorithms, making them easily customizable and ready for analysis. For our analysis, we employed MMYolo to first convert the weights from .pt (Pytorch model) to .pth (State dictionary file, i.e., weights, bias, etc.), and to second visualize the different activation maps of YOLOv8 during inference. MMYolo allows you to specify the model type, weight file, target layer, and channel reduction.
@@
YOLOv8 uses CSPDarknet53 [15] as its backbone [Figure 7], a deep neural network that extracts features at multiple resolutions (scales) by progressively down-sampling the input image. The feature maps produced at different resolutions contain information about objects at different scales in the image and different levels of detail and abstraction. YOLOv8 can incorporate different feature maps at different scales to learn about object shapes and textures, which helps it achieve high accuracy in most object detection tasks. YOLOv8’s backbone consists of four sections, each with a single convolution followed by a c2f module [19]. The c2f module is a new introduction to CSPDarknet53. The module comprises splits where one end goes through a bottleneck module (two 3x3 convolutions with residual connections). The bottleneck module output is further split N times, where N corresponds to the YOLOv8 model size. These splits are all finally concatenated and passed through one final convolution layer. This final layer is where we will get the activations.
@@
Figure 3 shows the original F-14 and F-18 images and the activations of the four c2f stages in the network, with each stage being more profound in the network from the second image right. The activation map corresponding to the shallowest c2f module shows the broadest activation. This module detects the two wings of the aircraft and determines that this object is a plane. The second activation map corresponds to the second c2f module in our backbone. It shows strong activations at different components of the aircraft, such as locating the wings, body, cockpit, and payload. It appears that this layer is attempting to infer what kind of aircraft is being presented in the image by highlighting these features. The third activation map is starting to dive into the individual textures of the components of the aircraft, presumably checking for minute differences in the jet’s structure. Finally, the model’s final c2f module activates extremely fine-grained details and outlines in the respective images. These similar feature activation maps could be the reason that the model confuses the two.
3 Results
3.1 Generalized Model
To highlight our results, we address three challenging conditions: (1) detecting and classifying extremely small objects, (2) identifying flying objects that blend into their background, and (3) classifying different types of flying objects. We examined the performance of our generalized model, [13], against these challenges. This is demonstrated in Figure 5, which features four images that represent the bird, drone, passenger airplane, and V22 classes.
The first of the four images showcases the model’s ability to identify distant birds. In the second image, the model was put to the test against a very small drone that occupied only 0.026% of the image size while also blending in with its background. The model still resulted in the correct detection and classification of the drone. The third image shows the model’s ability to identify a minute passenger airplane of size 0.063% of the image, which is also blended into its surroundings. Finally, the fourth image features a V22 aircraft, which is an underrepresented class and accounts for only 3.57% of the entire dataset. A V22 can easily be mistaken as a drone due to its vertical propeller positioning. Despite these two hindering characteristics and only taking up 0.14% of the entire image, the image exhibits the model’s ability to still identify the V22 with impressive accuracy, achieving a confidence score of 0.83.